How to Improve Resolution (e.g., to Resolve Closely Eluting or Co-eluting Compounds) in LC
Answer #1

The resolution of Chromatographic separation depends on -

R_{s}=\frac{N^{\frac{1}{2}}}{4}\times \frac{\alpha -1}{\alpha }\times \frac{k'}{k'+1}

So , to modify and improve resolution, we have 3 ways...

1st

Increase N where N is column efficiency . It can be done by increasing the number of theoretical plate...Also , using higher column length and smaller particle size, it can be done.

2nd

Increase \alpha where \alpha is selectivity . This can be done by choosing proper mobile phase .

3rd

Increase k'. This also can be done by changing the polarity strength of mobile phase.

By using different type of modifications , the resolution picture between two peaks look like-

to Start N and a Change in k remain the same kand a remain the same Increase in kand N Increase in remain the same